OpenStudy (anonymous):

First question! Need help asap. 1. Consider eight, eight-cubic centimeter (8 cm3) sugar cubes stacked so that they form a single 2 x 2 x 2 cube. How does the surface area of the single, large cube compare to the total surface area of the individual eight cubes? Report your answer as a ratio. Be sure to show all calculations leading to an answer.

OpenStudy (anonymous):

The 8cm3 is supposed to be 8cm^3

OpenStudy (michele_laino):

hint: the edge of a single cube is: L= third root(8)= 2cm and the total surface area is S= 6*L^2 = 6*2*2=...cm^2 the edge of the 2 x2 x2 cube is L_1= 2*2= 4cm and the total surface area is S_1= 6*L_1^2= 6* 4*4 =...cm^2 so, the requested ratio is: S_1/S =...?